OpenStudy (anonymous):

Can someone simplify this expression for me? -1x2-2x-4+2+2x-1x+1x2? for the 1x2, 2 is an exponent.

OpenStudy (anonymous):

someone please answer this thanx

OpenStudy (anonymous):

suppose x2 = a then your equation is -a-2x-4+2+2x=x+a

OpenStudy (anonymous):

that is - in the above equation, not =

OpenStudy (anonymous):

ugh... I mean can someone simplify the equation not add letters and stuff

OpenStudy (anonymous):

then first job is to add all the letters isn't it?

OpenStudy (anonymous):

nooooooooooooooo

OpenStudy (anonymous):

its x-2

OpenStudy (anonymous):

Bye! Thanks for completely not helping me. adios

OpenStudy (anonymous):

you refuse to follow the steps. its your problem dude, i can't do it for you. I can only show you how.

OpenStudy (anonymous):

okay "dude" anyone that is smart wouldnt say that

OpenStudy (anonymous):

and how would you know that?

OpenStudy (anonymous):

cuz i have a 4.0 gpa

OpenStudy (anonymous):

waaw!

OpenStudy (anonymous):

yeah im in harvard at the moment graduating soon

OpenStudy (anonymous):

great! doing what?

OpenStudy (anonymous):

I major in psycology and am gettig a teaching degree

OpenStudy (anonymous):

can you describe a situation that 55-7x would be used in

OpenStudy (anonymous):

I see, I thought they taught basic algebra in high school. Anyway, -1x2-2x-4+2+2x-1x+1x2 = -x^2 - 2x +4 +2 +2x -x +x^2. Now, you need to bunch up all the like terms together, so -x^2+x^2 -2x +2x -x -4 + 2 = -x -2 (because -x^2+x^2 = 0 and -2x+2x = 0 and -4 +2 = -2)

OpenStudy (anonymous):

-1x-2

OpenStudy (anonymous):

no they never when i went to high school, they may now tho

OpenStudy (anonymous):

I don't know, I didn't go to high school in the US.

OpenStudy (anonymous):

no do you see how i got -x -2 ?

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

so it is -1x-2?? Makes sense

OpenStudy (anonymous):

yes, x = 1x. you don't need to write the '1' explicitly.

OpenStudy (anonymous):

OK OK